Your firm is considering a project with a five-year life and an initial cost of $120,000. The discount rate for the project is 10%. The firm expects to sell 2,100 units a year. The expected cash flow per unit is $20. The firm will have the option to abandon this project after three years at which time it expects it could sell the project for $75,000. If this project sees units fall to 1,960 in the last two years, then which is the highest cash flow per unit that the firm will be willing to abandon?
Answer #1

NPV if abandon=-120000+2100*q/1.1+2100*q/1.1^2+2100*q/1.1^3+75000/1.1^3

NPV if doesnt abandon=-120000+2100*q/1.1+2100*q/1.1^2+2100*q/1.1^3+1960*q/1.1^4+1960*q/1.1^5

-120000+2100*q/1.1+2100*q/1.1^2+2100*q/1.1^3+1960*q/1.1^4+1960*q/1.1^5<=-120000+2100*q/1.1+2100*q/1.1^2+2100*q/1.1^3+75000/1.1^3

=>75000>=1960*q/1.1+1960*q/1.1^2

=>q<=75000/(1960/1.1+1960/1.1^2)
=>Cash flow per unit=22.04810496